The sum of three consecutive even integers is fifty-four. The sum of the first two consecutive integers is 34. What is the first consecutive integer?: The sum of three consecutive even integers is fifty-four. The sum of the first two consecutive integers is 34. What is the first consecutive integer?
Answer by nerdybill(1127)

The sum of three consecutive even integers is fifty-four. The sum of the first two consecutive integers is 34. What is the first consecutive integer?
.
Let x = 1st of three consecutive even integers
x+2 = 2nd of three consecutive even integers
x+4 = 3rd of three consecutive even integers
.
From: "The sum of three consecutive even integers is fifty-four." we get our equation:
x + (x+2) + (x+4) = 54
3x + 6 = 54
3x = 48
x = 16 (this is the 1st of three consecutive even integers)